# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v1

@@ Abstract.
where the implied constant is absolute and $N_{x}=P_{1}P_{2}\cdots$ is the unique factorization of $N_{x}$ into primes or ones with the $P_{i}$’s being non-increasing. This establishes a conjecture of Arratia (2002), who proved that the right-hand side in the above estimate can be made $\ll\log\!\log x$. We also use this coupling to give a probabilistic proof of the Dirichlet law for the average distribution of the integer factorization into $k$ parts proved in 2023 by Leung and we improve on its error term.
@@ 1. Introduction
is bounded by a positive monotone function that is $o(\log x)$ as $x\to\infty$. The random variables $N_{x}$ and $\mathbf{V}$ must be strongly correlated in this new probability space to achieve this. Indeed, if, for instance, $V_{1}$ and $P_{1}$ were independent, then we would have that $P_{1}\leqslant x^{1/3}$ and $V_{1}>2/3$ with positive probability, so that $\big{|}\log P_{1}-V_{1}\log x\big{|}\geqslant\frac{\log x}{3}$ with positive probability. Hence, a coupling with $N_{x}$ and $\mathbf{V}$ being independent (also called a trivial coupling) makes (1.1) $\gg\log x$. In 2002, Richard Arratia [2] constructed a coupling satisfying
@@
with $a(x)\coloneqq\frac{\log 2}{3\log x}$ for any coupling of $N_{x}$ and $\mathbf{V}$. However, this last expectation is exactly equal to $\log 2$ no matter how we choose $x\geqslant 2$.
@@
1.1. Application to distribution of factorizations of random integers
@@
Let $\omega(n)$ denote the number of prime factors of $n$. If $n$ is square-free and $f\in\mathcal{F}_{k}$, then properties (b) and (c) of the definition above imply that
@@
We will use the class of functions $\mathcal{F}_{k}(\bm{\alpha})$ to define certain “random factorizations” of a random integer into $k$ parts. Specifically, let us fix $f\in\mathcal{F}(\bm{\alpha})$ and $x\geqslant 1$. We then define the random $k$-factorization corresponding to $f$ to be a random variable $\mathbf{D}_{x}=(D_{x,1},\dots,D_{x,k})$ to be a random variable taking values on $\mathbb{N}^{k}$ and satisfies the formula [^2] Strictly speaking, we only need property (a) of Definition 1.1 to define $\mathbf{D}_{x}$. But we will also need the other two properties when proving Theorem 2 below.
@@
where $P_{1}P_{2}\cdots$ is the prime factorization of $N_{x}$ as before.
@@
$\mathbb{P}[D_{x}\leqslant N_{x}^{u}]=F_{(\alpha_{1},\alpha_{2})}(u)+O_{\alpha_{1},\alpha_{2}}\Big{(}(1+u\log x)^{-\alpha_{1}}(1+(1-u)\log x)^{-\alpha_{2}}\Big{)}.$
@@
We have a new approach to this problem: Since the size of the divisors on a logarithmic scale of any integer is entirely determined by the size of its prime factors, we might expect Leung’s theorem to be a consequence of some quantitative form of Billingsley’s Theorem. Indeed, using the coupling from Theorem 1, we deduce an improved form Leung’s theorem for the class of functions $\mathcal{F}_{k}(\bm{\alpha})$.
@@
The proof of Theorem 2 is based on a 1987 result of Donnelly and Tavaré [8], who proved the following probabilistic version of the Leung’s theorem: If $\mathbf{V}=(V_{1},V_{2},\ldots)$ is a Poisson–Dirichlet process and $(C_{i})_{i\geqslant 1}$ is a sequence of i.i.d. random variables supported on $\{1,\ldots,k\}$ satisfying $\mathbb{P}[C_{i}=j]=\alpha_{j}$, then
@@
When one of the $u_{i}$’s is close to 00 or $1$, the behavior changes completely in the error term of Theorem 2, because the vector $\bm{\delta}_{x}$ has discrete distribution (compared to the continuous distribution $\operatorname{Dir}(\bm{\alpha})$). For instance, if $u_{i}<\log 2/\log x$ for some $i$, then the relation $D_{x,i}\leqslant N_{x}^{u_{i}}$ implies that $D_{x,i}=1$.
@@ 1.2. Structure of the paper
We have organized the paper as follows in two main parts.
@@
  • In Section 8, we present an argument due to Donnelly–Tavaré showing that random $k$-partitions of the Poisson–Dirichlet distribution are distributed according to Dirichlet’s law.
  • In Section 9, we use the coupling of Section 2 to construct a coupling of the random $k$-factorization $\mathbf{D}_{x}$ and an analogous $k$-partition of the Poisson–Dirichlet distribution. We then use this coupling to reduce Theorem 2 to estimating certain boundary events, one involving number-theoretic objects and the other one pure probabilistic objects.
@@ 1.3. Notation
Moreover, $n\in\mathbb{N}$, we let $s(n)$ denote its large square-full divisor. Also, we let $n^{\flat}=n/s(n)$ and note that $n^{\flat}$ is square-free and co-prime to $s(n)$.
We write $\pi(x)$ for the number of primes $\leqslant x$. We shall also use heavily Chebyshev’s function $\theta(x)=\sum_{p\leqslant x}\log p$.
@@ Acknowledgements
The authors would like to thank Matilde Lalín and Gérald Tenenbaum for their helpful comments on the paper.
TH is supported by the Courtois Chair II in fundamental research.
DK is supported by the Courtois Chair II in fundamental research, by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2018-05699 and RGPIN-2024-05850) and by the Fonds de recherche du Québec - Nature et technologies (2022-PR-300951 and 2025-PR-345672).
@@ 2. The coupling
with $\gamma$ being the Euler-Mascheroni constant and $q_{j}=p_{j}^{v_{j}}$ being the $j$th smallest prime power, i.e., $(q_{j})_{j}$ is the sequence $2,3,2^{2},5,7,2^{3},3^{2},\dots$ Note that
@@
$\sum_{p^{k}\leqslant y}\frac{1}{kp^{k}}=\log\log y+\gamma+O(1/(\log y)^{3})\qquad(y\geqslant 2).$
@@
Indeed, the Prime Number Theorem (Proposition A.1) implies that $\psi(q_{j}+Cq_{j}/(\log q_{j})^{3})-\psi(q_{j})>0$ if $C$ is large enough, whence $q_{j+1}\leqslant q_{j}+Cq_{j}/(\log q_{i})^{3}$, as needed.
@@
    Construct the sequence of random prime powers or ones $(Q_{i})_{i\geqslant 1}$ by letting $Q_{i}\coloneqq e^{h(L_{i}\log x)}$. Note that only finitely many $Q_{i}$ will be prime powers.
@@
    Define the extra prime $P_{\mathtt{extra}}$ as the smallest element of the set $\{1\}\cup\{\text{primes}\}$ such that $\theta(P_{\mathtt{extra}})\geqslant U_{1}^{\prime}\theta(x/J_{x})$, where $\theta(y)=\sum_{p\leqslant y}\log p$ is Chebyshev’s function. (In particular, we have $P_{\mathtt{extra}}=1$ when $J_{x}>x/2$; otherwise, $P_{\mathtt{extra}}$ is a prime $\leqslant x/J_{x}$.)
@@ 3. The $\ell^{1}$ distance within the coupling
Finally, we note as well as that $\log P_{\mathtt{extra}}=\log M_{x}-\sum_{i\geqslant 2}\log Q_{i}$, $L_{1}=1-\sum_{i\geqslant 2}L_{i}$. Since we have also assumed that $M_{x}=N_{x}$, we have
@@
$\sum_{i\geqslant 2}|\log Q_{i}-L_{i}\log x|\leqslant\sum_{i\geqslant 2}r(L_{i}\log x)\leqslant\Theta_{x}.$
@@ 4. Another realization of the GEM distribution
Therefore, for the next sections, we will assume that the process $\mathbf{L}=(L_{1},L_{2},\ldots)$ sampled for our coupling was determined by $\mathscr{R}$ by defining $L_{1}\coloneqq 1-\frac{S_{1}}{\log x}$ and $L_{j}\coloneqq\frac{Y_{j-1}}{\log x}$ for $j\geqslant 2$. In this setting, we now have that
@@ 5. An integer-friendly version of $J_{x}$
Recall the Poisson Process $\mathscr{R}=(W_{i},Y_{i})_{i\in\mathbb{Z}}$ given in Definition 4.1. We then define
$\mathscr{R}^{*}\coloneqq\Big{\{}\Big{(}W_{i}Y_{i}/h(Y_{i}),e^{h(Y_{i})}\Big{)}:i\in\mathbb{Z},\ Y_{i}>e^{-\gamma}\Big{\}}.$
Without loss of generality, we may assume that the quantities $W_{i}Y_{i}/h(Y_{i})$ with $Y_{i}>e^{-\gamma}$ are all distinct. By the Mapping Theorem (Proposition B.3) applied to the map $(w,y)\to(wy/h(y),e^{h(y)})$, the random set $\mathscr{R}^{*}$ is a Poisson process on the space $\mathbb{R}_{>0}\times\{\text{prime powers}\}$ with mean measure $\mu^{*}$ satisfying
@@
for any $B\subseteq\mathbb{R}_{>0}$ and any $q\in\mathbb{N}$. There is a unique way to relabel the points of $\mathscr{R}^{*}$ as $\{(T^{*}_{i},Q^{*}_{i})\,:\,i\in\mathbb{Z}\}$ such that:
  • $T^{*}_{i}<T^{*}_{i+1}$ for all $i\in\mathbb{N}$;
  • $\prod_{i=1}^{\infty}Q_{i}^{*}\leqslant x<\prod_{i=0}^{\infty}Q_{i}^{*}$.
We then define the random integer
$J_{x}^{*}\coloneqq\prod_{i=1}^{\infty}Q^{*}_{i}.$
@@
$\mathbb{P}[Z=\ell]=(1-\lambda)\cdot\lambda^{\ell}\qquad\text{for}\ \ell=0,1,2\dots$
@@
$\mathbb{P}[Z=\ell]=\frac{1}{2\pi i}\oint_{|s|=1}\frac{\mathbb{E}[s^{Z}]}{s^{\ell+1}}\,\mathrm{d}s.\qed$
@@
For any $\ell\geqslant 1$, we have
$\mathbb{P}[Z=\ell]$ $=\sum_{\begin{subarray}{c}m_{1},\ldots,m_{\ell}\geqslant 0\\ \sum jm_{j}=\ell\end{subarray}}\Bigg{(}\prod_{j=1}^{\ell}\mathbb{P}[X_{j}=m_{j}]\prod_{j=\ell+1}^{\infty}\mathbb{P}[X_{j}=0]\Bigg{)}$
@@
whence $\mathbb{P}[Z=\ell]=(1-\lambda)\cdot\lambda^{\ell}$, as claimed. Finally, we compute $\mathbb{P}[Z=0]$ using the formula $\mathbb{P}[Z=0]=1-\mathbb{P}[Z\geqslant 1]$. ∎
@@
$\mathbb{P}[J_{x}^{*}=j]=\frac{1}{j\log x}\left(1+O\!\left(\frac{1}{\log x}\right)\right).$
@@
The family of random variables $(Z_{p}(t))_{p}$ is independent since $\mathscr{R}^{*}$ is a Poisson process. We also know that each random variable $N_{p^{k}}(t)$ follows a Poisson distribution with parameter $\frac{1}{kp^{k(1+t)}}$. Hence, $\mathbb{P}[Z_{p}(t)=\ell]=(1-p^{-1-t})p^{-\ell(1+t)}$ by Lemma 5.1 for all $\ell\geqslant 0$. Therefore,
$\mathbb{P}[I^{*}_{t}=j]=\prod_{p}\mathbb{P}[Z_{p}(t)=\nu_{p}(j)]=\frac{j^{-1-t}}{\zeta(1+t)}.$
All points of $\mathscr{R}^{*}$ have distinct $T^{*}$-coordinates with probability one. In this situation, we have that $J_{x}^{*}=j$ if and only if there exists exactly one point $(T^{*},Q^{*})\in\mathscr{R}^{*}$ such that $I^{*}_{T^{*}}=j$ and $Q^{*}>\frac{x}{j}$. Thus, we have
@@
(5.1) $\mathbb{P}[J_{x}^{*}=j]=\int_{0}^{\infty}\!\Bigg{(}\sum_{q>x/j}\frac{\Lambda(q)}{q^{1+t}}\Bigg{)}\cdot\frac{j^{-1-t}}{\zeta(1+t)}\,\mathrm{d}t.$

### v2

@@ Abstract.
where the implied constant is absolute and $N_{x}=P_{1}P_{2}\cdots$ is the unique factorization of $N_{x}$ into primes or ones with the $P_{i}$’s being non-increasing. This establishes a conjecture of Arratia (2002), who proved that the left-hand side in the above estimate can be made $\ll\log\!\log x$. We also use this coupling to give a probabilistic proof of the Dirichlet law for the average distribution of the integer factorization into $k$ parts proved in 2023 by Leung and we improve on its error term.
@@ 1. Introduction
is bounded by a positive monotone function that is $o(\log x)$ as $x\to\infty$. The random variables $N_{x}$ and $\mathbf{V}$ must be strongly correlated in this new probability space to achieve this. Indeed, if, for instance, $P_{1}$ and $V_{1}$ were independent, then we would have that $P_{1}\leqslant x^{1/3}$ and $V_{1}>2/3$ can happen at the same time with positive probability, which implies $\big{|}\log P_{1}-V_{1}\log x\big{|}\geqslant\frac{\log x}{3}$ with positive probability. Hence, a coupling with $N_{x}$ and $\mathbf{V}$ being independent (also called a trivial coupling) makes (1.1) $\gg\log x$. In 2002, Richard Arratia [2] constructed a coupling satisfying
@@
with $a(x)\coloneqq\frac{\log 2}{3\log x}$, no matter how we choose the coupling between $N_{x}$ and $\mathbf{V}$. However, the expectation above is exactly equal to $\log 2$ for any choice of $x\geqslant 2$.
@@
1.1. Application to the distribution of factorizations of random integers
@@
Let $\omega(n)$ denote the number of distinct prime factors of $n$. If $n$ is square-free and $f\in\mathcal{F}_{k}(\bm{\alpha})$, then properties (b) and (c) of the definition above imply that
@@
We will use the class of functions $\mathcal{F}_{k}(\bm{\alpha})$ to define certain “random factorizations” into $k$ parts of a random integer. Specifically, let us fix $f\in\mathcal{F}_{k}(\bm{\alpha})$ and $x\geqslant 1$. We then define a random $k$-factorization corresponding to $f$ to be a random vector $\mathbf{D}_{x}=(D_{x,1},\dots,D_{x,k})$ taking values on $\mathbb{N}^{k}$ and satisfies the formula [^2] Strictly speaking, we only need property (a) of Definition 1.1 to define $\mathbf{D}_{x}$. But we will also need the other two properties when proving Theorem 2 below.
@@
where $P_{1}P_{2}\cdots$ is the prime factorization of $N_{x}$ as before. With this definition, the vectors $(\nu_{p}(D_{x,1}),\cdots,\nu_{p}(D_{x,k}))$ all follow multinomial distributions for every prime $p$.
@@
$\mathbb{P}\big{[}D_{x}\leqslant N_{x}^{u}\big{]}=F_{(\alpha_{1},\alpha_{2})}(u)+O_{\alpha_{1},\alpha_{2}}\Big{(}(1+u\log x)^{-\alpha_{1}}(1+(1-u)\log x)^{-\alpha_{2}}\Big{)}.$
@@
We have a new approach to this problem: Since the size of the divisors on a logarithmic scale of any integer is entirely determined by the size of its prime factors, we might expect Leung’s theorem to be a consequence of some quantitative form of Billingsley’s Theorem. Indeed, using the coupling from Theorem 1, we deduce an improved form of Leung’s theorem for the class of functions $\mathcal{F}_{k}(\bm{\alpha})$.
@@
The proof of Theorem 2 is based on a 1987 result of Donnelly and Tavaré [8], who proved the following probabilistic version of Leung’s theorem: If $\mathbf{V}=(V_{1},V_{2},\ldots)$ is a Poisson–Dirichlet process and $(C_{i})_{i\geqslant 1}$ is a sequence of i.i.d. random variables supported on $\{1,\ldots,k\}$ satisfying $\mathbb{P}[C_{i}=j]=\alpha_{j}$, then
@@
When one of the $u_{i}$’s is close to 00 or $1$, the behavior changes completely in the error term of Theorem 2, because the vector $\bm{\delta}_{x}$ has a discrete distribution (compared to the continuous distribution $\operatorname{Dir}(\bm{\alpha})$). For instance, if $u_{i}<\log 2/\log x$ for some $i$, then the relation $D_{x,i}\leqslant N_{x}^{u_{i}}$ implies that $D_{x,i}=1$.
@@ 1.2. Structure of the paper
We have organized the paper in two main parts.
@@
  • In Section 8, we present the argument due to Donnelly–Tavaré showing that random $k$-partitions of the components of a Poisson–Dirichlet process are distributed according to a Dirichlet law.
  • In Section 9, we use the coupling of Section 2 to construct a coupling of the random $k$-factorization $\mathbf{D}_{x}$ and an analogous $k$-partition of the Poisson–Dirichlet distribution. We then use this coupling to reduce Theorem 2 to estimating two boundary events, one involving number-theoretic objects and the other one using pure probabilistic objects.
@@ 1.3. Notation
Moreover, for $n\in\mathbb{N}$, we let $s(n)$ denote its largest square-full divisor. Also, we let $n^{\flat}\coloneqq n/s(n)$ and we note that $n^{\flat}$ is square-free and co-prime to $s(n)$.
We write $\pi(x)$ for the number of primes $\leqslant x$. We shall also use heavily Chebyshev’s function $\theta(x)\coloneqq\sum_{p\leqslant x}\log p$.
@@ Acknowledgements
The authors would like to thank Matilde Lalín, Sun-Kai Leung and Gérald Tenenbaum for their helpful comments on the paper.
TH is supported by the Courtois Chair II in fundamental research. DK is supported by the Courtois Chair II in fundamental research, by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2018-05699 and RGPIN-2024-05850) and by the Fonds de recherche du Québec - Nature et technologies (2022-PR-300951 and 2025-PR-345672).
@@ 2. The coupling
with $\gamma$ being the Euler-Mascheroni constant and $q_{j}=p_{j}^{v_{j}}$ being the $j$th smallest prime power, i.e., $(q_{j})_{j\geqslant 1}$ is the sequence $2,3,2^{2},5,7,2^{3},3^{2},\dots$ Note that
@@
$\sum_{p^{k}\leqslant y}\frac{1}{kp^{k}}=\log_{2}y+\gamma+O(1/(\log y)^{3})\qquad(y\geqslant 2).$
@@
Indeed, the Prime Number Theorem (Proposition A.1) implies that $\theta(q_{j}+Cq_{j}/(\log q_{j})^{3})-\theta(q_{j})>0$ if $C$ is large enough, whence $q_{j+1}\leqslant q_{j}+Cq_{j}/(\log q_{j})^{3}$, as needed.
@@
    Construct the sequence of random prime powers or ones $(Q_{i})_{i\geqslant 1}$ by letting $Q_{i}\coloneqq e^{h(L_{i}\log x)}$. Note that $Q_{i}=1$ whenever $L_{i}\leqslant\frac{e^{-\gamma}}{\log x}$ so there are only finitely many $Q_{i}$’s that are prime powers.
@@
    Define the extra prime $P_{\mathtt{extra}}$ as the smallest element in the set $\{1\}\cup\{\text{primes}\}$ that would satisfy $\theta(P_{\mathtt{extra}})\geqslant U_{1}^{\prime}\theta(x/J_{x})$, where $\theta(y)=\sum_{p\leqslant y}\log p$ is Chebyshev’s function. (In particular, we have $P_{\mathtt{extra}}=1$ when $J_{x}>x/2$; otherwise, $P_{\mathtt{extra}}$ is a prime $\leqslant x/J_{x}$.)
@@ 3. The $\ell^{1}$ distance within the coupling
Finally, we also note that $\log P_{\mathtt{extra}}=\log M_{x}-\sum_{i\geqslant 2}\log Q_{i}$ and $L_{1}=1-\sum_{i\geqslant 2}L_{i}$. Since we have assumed that $M_{x}=N_{x}$, we have
@@
$\sum_{i\geqslant 2}|\log Q_{i}-L_{i}\log x|=\sum_{i\geqslant 2}r(L_{i}\log x)\leqslant\Theta_{x}.$
@@ 4. Another realization of the GEM distribution
For the next sections and the proof below, we will assume that the process $\mathbf{L}=(L_{1},L_{2},\ldots)$ sampled for our coupling was determined by $\mathscr{R}$ by defining $L_{1}\coloneqq 1-\frac{S_{1}}{\log x}$ and $L_{j}\coloneqq\frac{Y_{j-1}}{\log x}$ for $j\geqslant 2$. In this setting, we now have that
@@ 5. An integer-friendly version of $J_{x}$
Recall the Poisson Process $\mathscr{R}$ given in Definition 4.1. We then define
$\mathscr{R}^{*}\coloneqq\Big{\{}\Big{(}WY/h(Y),e^{h(Y)}\Big{)}:(W,Y)\in\mathscr{R},\ Y>e^{-\gamma}\Big{\}}.$
Without loss of generality, we may assume that the quantities $WY/h(Y)$ with $(W,Y)\in\mathscr{R}$ and $Y>e^{-\gamma}$ are all distinct. By the Mapping Theorem (Proposition B.3) applied to the map $(w,y)\mapsto(wy/h(y),e^{h(y)})$ on the Poisson process $\mathscr{R}$ restricted to $\mathbb{R}_{>0}\times\mathbb{R}_{>e^{-\gamma}}$, the random set $\mathscr{R}^{*}$ is a Poisson process on the space $\mathbb{R}_{>0}\times\{\text{prime powers}\}$ with mean measure $\mu^{*}$ satisfying
@@
for any $B\subseteq\mathbb{R}_{>0}$ and any $q\in\mathbb{N}$. Note that the $w$-coordinates of the points in $\mathscr{R}$ restricted to $\mathbb{R}_{>0}\times\mathbb{R}_{>e^{-\gamma}}$ are now bounded with probability one. For each such realization of the points of $\mathscr{R}^{*}$, there is a unique labeling of them as $\{(T^{*}_{i},Q^{*}_{i})\,:\,i\in\mathbb{Z}_{\leqslant K}\}$ in such a way that the following properties hold:
  • $T^{*}_{i-1}<T^{*}_{i}$ for all $i\in\mathbb{Z}_{\leqslant K}$;
  • $\prod_{i=1}^{K}Q_{i}^{*}\leqslant x<\prod_{i=0}^{K}Q_{i}^{*}$.
(Note that $K$ is a random variable and is not fixed.) We then define the random integer
$J_{x}^{*}\coloneqq\prod_{i=1}^{K}Q^{*}_{i}.$
@@
$\mathbb{P}\big{[}Z=\ell\big{]}=(1-\lambda)\cdot\lambda^{\ell}\qquad\text{for}\ \ell=0,1,2\dots$
@@
$\mathbb{P}\big{[}Z=\ell\big{]}=\frac{1}{2\pi i}\oint_{|s|=1}\frac{\mathbb{E}[s^{Z}]}{s^{\ell+1}}\,\mathrm{d}s.\qed$
@@
For the case $\ell\geqslant 1$, we have
$\mathbb{P}\big{[}Z=\ell\big{]}$ $=\sum_{\begin{subarray}{c}m_{1},\ldots,m_{\ell}\geqslant 0\\ \sum jm_{j}=\ell\end{subarray}}\Bigg{(}\prod_{j=1}^{\ell}\mathbb{P}[X_{j}=m_{j}]\prod_{j=\ell+1}^{\infty}\mathbb{P}[X_{j}=0]\Bigg{)}$
@@
whence $\mathbb{P}[Z=\ell]=(1-\lambda)\cdot\lambda^{\ell}$, as claimed. Finally, for the case $\ell=0$, we recover $\mathbb{P}[Z=0]$ using the formula $\mathbb{P}[Z=0]=1-\mathbb{P}[Z\geqslant 1]$. ∎
@@
$\mathbb{P}\big{[}J_{x}^{*}=j\big{]}=\frac{1}{j\log x}\left(1+O\!\left(\frac{1}{\log x}\right)\right).$
@@
The family of random variables $(Z_{p}(t))_{p}$ is independent since $\mathscr{R}^{*}$ is a Poisson process. We also know that each random variable $N_{p^{k}}(t)$ follows a Poisson distribution with parameter $\frac{1}{kp^{k(1+t)}}$. Hence, $\mathbb{P}\big{[}Z_{p}(t)=\ell\big{]}=(1-p^{-1-t})p^{-\ell(1+t)}$ by Lemma 5.1 for all $\ell\geqslant 0$, and we have
$\mathbb{P}\big{[}I^{*}_{t}=j\big{]}=\prod_{p}\mathbb{P}\big{[}Z_{p}(t)=\nu_{p}(j)\big{]}=\frac{j^{-1-t}}{\zeta(1+t)}$
for all $j\geqslant 1$. All points of $\mathscr{R}^{*}$ have distinct $T^{*}$-coordinates with probability one. In this situation, we have that $J_{x}^{*}=j$ if, and only if, there exists exactly one point $(T^{*},Q^{*})\in\mathscr{R}^{*}$ such that $I^{*}_{T^{*}}=j$ and $Q^{*}>\frac{x}{j}$. Thus, we have
@@
(5.1) $\mathbb{P}\big{[}J_{x}^{*}=j\big{]}=\int_{0}^{\infty}\!\Bigg{(}\sum_{q>x/j}\frac{\Lambda(q)}{q^{1+t}}\Bigg{)}\cdot\frac{j^{-1-t}}{\zeta(1+t)}\,\mathrm{d}t.$